OpenStudy (anonymous):

A triangle has a base 14 cm long and an altitude 8 cm long. What is the area of the triangle?

OpenStudy (anonymous):

Area = 1/2 * base * heigth Area = 1/2 * 14 * 8 cm^2

OpenStudy (anonymous):

A. 22 cm2 B. 44 cm2 C. 56 cm2 D. 112 cm2

OpenStudy (anonymous):

You can't computer 1/2 * 14 * 8?

OpenStudy (anonymous):

Just use the area = (1/2) * base * height formula. What do you get?

OpenStudy (anonymous):

C is the answer.